Calculate the present worth of all costs for a newly acquired machine with an initial cost of $34,000, no trade-in value, a life of 14 years, and an annual operating cost of $17,000 for the first 3 years, increasing by 10% per year thereafter. Use an interest rate of 10% per year.

The present worth of all costs for a newly acquired machine is determined to be $
Answer #1

initial investment = 34000

O&M cost = 17000 for first three years and 10% increase after that (g=10%)

useful life = 14 yrs

i = 10%

The present value of gradient series when i=g is given by formula = C*n/(i+1)

Here n =12 (14-2) becaue gradient series is from yr 3 to yr 14

we will consider gradient series from yr 3 to yr 14 and present value at the end of second of this gradient series

= -17000 * 12 / (1+0.1) = 17000 * 12 / 1.1

= -185454.54

Now present value of maintenace cost = -17000/1.1 - 17000/1.12 - 185454.54/1.12

= -182772.347

Net present value of cost = -34000-182772.347 = -216772.34

We could have used excel and gotten the same result, using excel

Year Investment O&M cost Net cash flow discount factor Present value
0 -34000 -34000 1.000000 -34000.00
1 -17,000.00 -17,000.00 0.909091 -15454.55
2 -17,000.00 -17,000.00 0.826446 -14049.59
3 -17,000.00 -17,000.00 0.751315 -12772.35
4 -18,700.00 -18,700.00 0.683013 -12772.35
5 -20,570.00 -20,570.00 0.620921 -12772.35
6 -22,627.00 -22,627.00 0.564474 -12772.35
7 -24,889.70 -24,889.70 0.513158 -12772.35
8 -27,378.67 -27,378.67 0.466507 -12772.35
9 -30,116.54 -30,116.54 0.424098 -12772.35
10 -33,128.19 -33,128.19 0.385543 -12772.35
11 -36,441.01 -36,441.01 0.350494 -12772.35
12 -40,085.11 -40,085.11 0.318631 -12772.35
13 -44,093.62 -44,093.62 0.289664 -12772.35
14 -48,502.98 -48,502.98 0.263331 -12772.35
Net present value -216772.35